The chemistry lab at the University of Hardwoods keeps two acidic solutions on hand. One is 20% acid and the other is 35% acid. How much 20% acid solution and how much 35% acid solution should be used to prepare 25 liters of a 26% acid solution?
Answer by stanbon(75887)

One is 20% acid and the other is 35% acid. How much 20% acid solution and how much 35% acid solution should be used to prepare 25 liters of a 26% acid solution?
----------------
Quantity Eq: x + y = 25 liters
----
Acid Eq::0.20x + 0.35y = 0.26*25
--------------------------------------
Multiply the Quantity Eq. by 20
Multiply the Acid Eq. by 100
-------------------------------------
20x + 20y = 20*25
20x + 35y = 26*25
----
Subtract and solve for "y":
15y = 6*25
y = 2*5 = 10 liters (amt. of 35% solution needed in the mixture)
---
Siince x+y = 25, x = 15 liters (amt. of 20% solution needed in the mixture)
========================
Cheers,
Stan H.